OpenStudy (anonymous):

A railroad tunnel is shaped like a semi ellipse. The height of the tunnel at the center is 54 ft tall and the vertical clearance must be 18 ft tall at a point 8 ft from the center. Find an equation for the ellipse.

OpenStudy (anonymous):

Equation of an ellipse (x - h)^2 / a^2 + (y - k)^2 / b^2 = 1 (h , k) is the centre of the ellipse a is the radius in the x-axis direction and b is the radius in the y-axis direction

OpenStudy (anonymous):

SO would the answer be x^2/8^2+y^2/18^2=1?

OpenStudy (anonymous):

I don't really know because I don't understand the question.

OpenStudy (anonymous):

Okay. Thanks for the equation though.